OpenStudy (anonymous):

Another implicit differential question. Regard y as the independent variable and x as the dependent variable and use implicit differentiation to find dx/dy. (x^3)(y^2)-(x^3)y+2x(y^3)=0

OpenStudy (anonymous):

here is what i did:

OpenStudy (anonymous):

\[3x^2y^2\frac{ dx }{ dy }-(3x^2y \frac{ dx }{ dy }+x^3)+(2(y^3\frac{ dx }{ dy }+3xy^2))\] \[3x^2y^2\frac{ dx }{ dy }-3x^2y \frac{ dx }{ dy }+2y^3\frac{ dx }{ dy }=x^3-2y^3-6xy^2\] \[\frac{ dx }{ dy }=\frac{ (x^3-2y^3-6x^2y) }{ 2y^3 }\]

OpenStudy (anonymous):

but apparently this is wrong, can anyone else find my mistake

zepdrix (zepdrix):

Ooo I see :O You forgot to product rule on the first term.

zepdrix (zepdrix):

maybe? hold on checking again XD

OpenStudy (anonymous):

my bad just forgot to put in this, but on paper i did put it

zepdrix (zepdrix):

\[\large\rm x^3y^2-x^3y+2xy^3=0\]It's somewhere in the derivative of the last term. Hmm.

OpenStudy (anonymous):

ooh i know what i messed up in

OpenStudy (anonymous):

never mind, its still wrong lol

zepdrix (zepdrix):

This is another sneaky trick you could have applied to the left side, \[\large\rm x^3(y^2-y)+2xy^3=0\]which saves you from doing the product rule one extra time. Just a thought :)

OpenStudy (anonymous):

I don't understand what you did, can you elaborate?

zepdrix (zepdrix):

I factored an x^3 out of the first two terms: \(\large\rm \color{orangered}{x^3}y^2-\color{orangered}{x^3}y=\color{orangered}{x^3}(y^2-y)\) Then you do your product rule a little bit differently for that first big term, \[\large\rm \color{royalblue}{(x^3)'}(y^2-y)+x^3\color{royalblue}{(y^2-y)'}+\color{royalblue}{(2x)'}y^3+2x\color{royalblue}{(y^3)'}=0\]

OpenStudy (anonymous):

oh ok

zepdrix (zepdrix):

It allows you to do product rule twice, instead of three times. Not a big deal though.

zepdrix (zepdrix):

And I recommend using \(\large\rm x'\) instead of \(\large\rm \frac{dx}{dy}\). It makes the math so much easier to read. But that's just my own preference I guess :)

zepdrix (zepdrix):

I can't seem to figure out where this is coming from:\[\large\rm 3x^2y^2\frac{ dx }{ dy }-3x^2y \frac{ dx }{ dy }+2y^3\frac{ dx }{ dy }=x^3\color{red}{-2y^3}-6xy^2\]The one in red. Hmm

OpenStudy (anonymous):

what about it?

OpenStudy (anonymous):

thats from the first term

OpenStudy (anonymous):

oh wait hmm

OpenStudy (anonymous):

ooh thats suppose to 2yx^3

OpenStudy (anonymous):

I must have messed when I was collecting like terms

OpenStudy (amistre64):

(x^3)(y^2) -(x^3)y +2x(y^3) =0 (x^3)'(y^2)+(x^3)(y^2)' -(x^3)'y-(x^3)y' +2x'(y^3)+2x(y^3)' =0'

OpenStudy (anonymous):

Can you guys check if this is right? \[\frac{ x^3-2yx^3-6xy^2 }{ 3x^2 y^2-3x^2y+2y^3}\]

OpenStudy (anonymous):

this is what I got for my final answer after fixing some mistakes

zepdrix (zepdrix):

Ooo that looks better! :)

OpenStudy (amistre64):

(3x^2x')(y^2)+(x^3)(2yy') -(3x^2x')y-(x^3)y' +2x'(y^3)+2x(3y^2 y') =0 y' = dy/dy = 1 (3x^2x')(y^2)+(x^3)(2y) -(3x^2x')y-(x^3) +2x'(y^3)+2x(3y^2) =0 x'* (3x^2)(y^2) -(3x^2)y +2(y^3) = -2x(3y^2)+(x^3)-(x^3)(2y) etc ...

OpenStudy (anonymous):

awsome thanks guys